# Chan-wei Hu1, Chengxuan Qian1, Huaxiu Yao4, Zhengzhong Tu [^1] footnotemark: 1

### v2

@@ Abstract
Shuo Xing1thanks:  Email: {shuoxing,tzz}@tamu.edu, Yuping Wang2, Peiran Li1, Ruizheng Bai1, Yueqi Wang3, Chan-wei Hu1, Chengxuan Qian1, Huaxiu Yao4, Zhengzhong Tu [^1] footnotemark: 1thanks:   Corresponding author. 1Texas A&M University  2University of Michigan  3UIUC  4UNC Chapel Hill
@@ 1 Introduction
The recent emergence of powerful Vision Language Models (VLMs) (Li et al., 2022, 2023a; Liu et al., 2024a; Li et al., 2024b; Meta, 2024; Bai et al., 2023; Wang et al., 2024b; Lu et al., 2024; Wu et al., 2024) has significantly extended the capabilities of Large Language Models (LLMs) (Devlin et al., 2018; Radford et al., 2019; Brown et al., 2020; Team et al., 2023; Roziere et al., 2023; Touvron et al., 2023a, b; Raffel et al., 2020; Yang et al., 2024; Team, 2024) into the visual domain, paving the way for innovative real-world applications that integrate multimodal information (Moor et al., 2023; Li et al., 2024a; Shao et al., 2024; Xing et al., 2024b; Rana et al., 2023; Kim et al., 2024). Despite their promising performance, VLMs remain susceptible to hallucinations—instances where the model produces outputs containing inaccurate or fabricated details about objects, attributes, and the logical relationships inherent in the input image (Rohrbach et al., 2018; Bai et al., 2024). Several factors contribute to this cross-modal inconsistency, including the separate low-quality or biased training data, imbalanced model architectures, and the disjoint pretraining of the vision encoder and LLM-backbone (Cui et al., 2023; Bai et al., 2024; Zhou et al., 2024a).
To mitigate the hallucinations in VLMs, the Directed Preference Optimization (DPO) techniques have been widely adopted (Deng et al., 2024; Zhou et al., 2024a; Fang et al., 2024; Zhou et al., 2024b; Guo et al., 2024; Chen et al., 2024; Wang et al., 2024c; Yu et al., 2024b; Li et al., 2023b; Wang et al., 2024a; Xiao et al., 2025; Xie et al., 2024; Fu et al., 2024). This involves constructing datasets enriched with human preference signals specifically targeting hallucinations, and then finetuning the models using algorithms like Direct Preference Optimization (DPO) (Rafailov et al., 2024). Existing methods generate the preference data by perturbing the ground truth responses (Zhou et al., 2024a) and corrupting the visual inputs/embeddings (Deng et al., 2024; Amirloo et al., 2024) to generate rejected responses or correcting/refining responses to produce chosen responses (Chen et al., 2024; Yu et al., 2023a). While methods based on response refinement yield the most reliable preference signals, they face scalability challenges due to the significant costs of manual correction processes. Conversely, directly corrupting input visual information or ground truth responses is overly simplistic, as this brute-force approach fails to generate plausible and natural hallucinations in a controlled manner. Moreover, during fine-tuning, directly applying DPO may cause the model to overly prioritize language-specific preferences, which potentially leads to suboptimal performance and an increased propensity for hallucinations (Wang et al., 2024a).
@@ 2 Preliminaries
Image retrieval aims to find relevant images from large databases – such as vector databases or indexed corpora – based on semantic similarity criteria. In this paper, we convert all images into vector representations and utilize the cosine similarity metric to evaluate their proximity to a reference image. The similarity between two images, $v_{1}$ and $v_{2}$, is computed as follows:
$s=\bigg{<}\frac{f_{p}(v_{1})}{||f_{p}(v_{1})||},\frac{f_{p}(v_{2})}{||f_{p}(v_{2})||}\bigg{>},$
where <,>$\cdot$,$\cdot$>< ⋅ , ⋅ > denotes the inner product in $l_{2}$ space, $f_{p}(v_{i})$ represents the image embeddings generated by the vision encoder $f_{v}(\cdot)$ of VLMs. In this paper, we employ the FAISS library (Douze et al., 2024; Johnson et al., 2019) for efficient vector searches, retrieving the top-$k$ most relevant images.
@@ 3 Methods
As shown in Figure 2, the process begins with an advanced VLM generating chosen responses from input images from the training set. A selective masking process is then applied, strategically omitting segments associated with objects, attributes, or logical relationships identified in the image. Next, leveraging the retrieved image from the same training dataset and the masked responses, the hallucination-prone VLM is prompted to complete the masked elements, obtaining rejected responses. The generated preference pairs (chosen vs. rejected) are then used to fine-tune the VLM with $\mathcal{L}_{\text{rDPO}}$ (eq. 1), a preference objective that integrates both visual and textual information to penalize hallucinations and reinforce grounded reasoning. Algorithm 1 in Appendix A provides an overview of Re-Align, while the detailed process is explained in the following subsections.
@@ 3.1 Preference Generation
Generating high-quality preference data, which includes both accurate ground-truth responses and controlled hallucinated examples, is crucial for effective preference optimization in pre-trained VLMs. Existing methods construct preference data by perturbing ground-truth responses (Zhou et al., 2024a), corrupting visual inputs/embeddings (Deng et al., 2024; Amirloo et al., 2024) to create rejected responses, or refining responses to obtain chosen responses (Chen et al., 2024; Yu et al., 2023a). Refinement produces high-quality preference data but comes at a high cost, whereas direct corruption is more scalable yet tends to generate unrealistic hallucinations and fails to produce plausible, natural ones in a controlled manner. To address these limitations, we introduce a novel image retrieval-based pipeline for preference data construction as shown in Figure 3, which consists of three key stages:
@@ 3.2 Preference Optimization
Methods SQA TextVQA MM-Vet VisWiz LLaVABench MMEP MMEC MMBench Avg. Rank
@@ 4.1 Re-Align for VLMs Alignment
We conducted experiments on both hallucination detection and general VQA tasks. Specifically, we assess our method’s performance in hallucination detection using the POPE dataset (Li et al., 2023c) and HallusionBench (Guan et al., 2023). For general VQA tasks, we leverage a diverse suite of benchmarks including ScienceQA (Lu et al., 2022), TextVQA (Singh et al., 2019), MM-Vet (Yu et al., 2023b), VisWiz (Gurari et al., 2018), LLaVABench (Liu, 2023), MME (Fu et al., 2023), and MMBench (Liu et al., 2024c).
@@
We sample 11k images from the LLaVA-Instruct-150K dataset (Liu et al., 2024a) to construct preference data, as illustrated in Figure 3. These images are initially used to generate QA pairs based on image captions and simple VQA tasks using GPT-4o mini (OpenAI, 2024). Furthermore, the images are encoded using clip-vit-large-patch14 (Radford et al., 2021a) to construct the knowledge base for image retrieval. For rejected responses, we use GPT-4o mini to mask the chosen response and all-mpnet-base-v2 (Reimers and Gurevych, 2019) to compute the similarity between the completed masked response and the original chosen response. We use LLaVA-v1.5-7B (Liu et al., 2024a) and LLaVA-v1.6-Mistral-7B (Li et al., 2024b) as our backbone models and perform Re-Align finetuning for $1$ epoch. All evaluations are conducted with a temperature setting of 0, and baseline results are reproduced using open-sourced model weights.
@@ 4.2 Scalability and Generalizability
  • Image-to-Text models: the typical architecture of VLMs, where a vision encoder is integrated with an LLM to enable cross-modal understanding. In this section, we evaluate Re-Align on LLaVA-v1.5-7B (Liu et al., 2024a), LLaVA-v1.5-13B (Liu et al., 2024a), LLaVA-v1.6-Vicuna-7B (Li et al., 2024b), and LLaVA-v1.6-Vicuna-13B (Li et al., 2024b).
@@ Results
Table 3 presents the performance of Re-Align using both standard image-to-text and unified VLM backbones across model sizes from 1B to 13B on the POPE benchmark (Li et al., 2023c). In experiments with the LLaVA-v1.5 series (Liu et al., 2024a), none of the baseline approaches consistently improve performance for either the 7B or the 13B models, highlighting the limited scalability of these methods. In contrast, Re-Align achieved substantial performance gains, outperforming both the baseline models and the vanilla version—most notably on the LLaVA-v1.5-13B variant. Similarly, experiments with the LLaVA-v1.6-Vicuna series (Li et al., 2024b) revealed the same trend, further underscoring Re-Align’s superior scalability. For unified vision-language models, especially Janus-Pro, integrating Re-Align yields a significant performance boost. Notably, Janus-Pro-1B experiences the greatest improvement, underscoring Re-Align’s robustness across different model architectures. However, Janus-Pro-1B, being the smallest among the evaluated VLMs, also exhibits the poorest overall performance on POPE, suggesting a correlation between model size and the propensity for hallucinations.
@@ 4.3 Ablation Study
In this section, we conduct a comprehensive ablation study to explore how the data curation framework and design of the objective function affect the Re-Align’ performance. The experimental setup follows the same setting as VLMs alignment experiments.
@@
Table 4: Impact of similarity threshold $\tau$ for generating the rejected responses in Re-Align across general and hallucination benchmarks for VLMs, and comparisons with baselines.
@@
In data curation, we generate preference data by inducing hallucinations at the segment level. To further investigate the impact of finer-grained perturbations, we conduct experiments using sentence-level masking. As shown in Table 5, using a sentence-level masking strategy, Re-Align still demonstrates significant improvement in reducing hallucination in VLMs. However, this approach leads to a slight drop in performance on general VQA tasks. More discussions on the masking strategy can be found in Appendix F.
Masking Strategy SQA TextVQA POPEr POPEp POPEa
@@
Table 5: Impact of masking strategy across general and hallucination benchmarks for VLMs, and comparisons with baselines.
@@
Results
Table 6 summarizes the performance of Re-Align when using both standard DPO and rDPO as the direct optimization objectives, evaluated on general VQA and hallucination tasks with LLaVA-v1.5-7B (Liu et al., 2024a) and LLaVA-v1.6-Mistral-7B (Li et al., 2024b) as backbones. The results indicate that employing rDPO as the finetuning objective consistently yields superior performance over standard DPO across both task categories, highlighting the benefits of incorporating visual preference signals during the alignment process for VLMs. Notably, even when solely employing DPO, Re-Align not only achieves performance gains over the vanilla models but also outperforms the baselines evaluated in the VLM alignment experiments on several tasks. This underscores the effectiveness of our image retrieval-based preference data construction.
@@ 6 Related Work
Reinforcement Learning from Human Feedback (RLHF) has emerged as a crucial technique for incorporating human preference signals into machine learning methods and models (Dong et al., 2024). RLHF frameworks can be broadly categorized into deep RL-based approaches and direct preference learning approaches. In deep RL-based methods, a reward model is first constructed, after which Proximal Policy Optimization (PPO) (Schulman et al., 2017; Christiano et al., 2017; Ziegler et al., 2019) is employed to optimize the reward signals with KL regularization (Ouyang et al., 2022; Touvron et al., 2023b). While the direct preference learning approaches optimize a designed loss target on the offline preference dataset directly, eliminating the need for a separate reward model(Rafailov et al., 2024; Azar et al., 2024; Tang et al., 2024; Ethayarajh et al., 2024).
@@
Large Vision Language Models (VLMs) (Li et al., 2022, 2023a; Liu et al., 2024a; Li et al., 2024b; Meta, 2024; Bai et al., 2023; Wang et al., 2024b; Lu et al., 2024; Wu et al., 2024) extended the understanding and reasoning capabilities of Large Language Models (LLMs) (Devlin et al., 2018; Radford et al., 2019; Brown et al., 2020; Team et al., 2023; Roziere et al., 2023; Touvron et al., 2023a, b; Raffel et al., 2020; Yang et al., 2024; Team, 2024; Pan et al., 2023) into the visual domain. By integrating vision encoders, such as CLIP (Radford et al., 2021b), image patches are first converted into embeddings and then projected to align with text embedding space, unlocking unprecedented cross-modal applications in the real world, such as biomedical imaging (Moor et al., 2023; Li et al., 2024a), autonomous systems (Shao et al., 2024; Tian et al., 2024; Sima et al., 2023; Xing et al., 2024b; Ma et al., 2025; Wang et al., 2025), and robotics (Rana et al., 2023; Kim et al., 2024; Xing et al., 2025).
@@
Current VLMs often suffer from hallucinations, producing inaccurate or misleading information that fails to accurately represent the content of the provided image (Zhu et al., 2024; Bai et al., 2024; Qian et al., 2025). Such misalignments can have catastrophic consequences when these models are deployed in real-world scenarios (Xing et al., 2024a). To address cross-modality hallucinations, recent research has primarily focused on applying direct preference optimization (Deng et al., 2024; Zhou et al., 2024a; Fang et al., 2024; Zhou et al., 2024b; Guo et al., 2024; Chen et al., 2024; Wang et al., 2024c; Yu et al., 2024b; Li et al., 2023b; Wang et al., 2024a) or contrastive learning (Sarkar et al., 2024) on the curated datasets with preference signals, and utilizing model editing techniques (Liu et al., 2024b; Yu et al., 2024a).

### v3

# Chan-Wei Hu1,  Chengxuan Qian1,  Huaxiu Yao4,  Zhengzhong Tu1

@@ Abstract
Shuo Xing1,  Peiran Li1,  Yuping Wang2,  Ruizheng Bai1,  Yueqi Wang3, Chan-Wei Hu1,  Chengxuan Qian1,  Huaxiu Yao4,  Zhengzhong Tu1thanks: Corresponding author. 1Texas A&M University  2University of Michigan  3UIUC  4UNC Chapel Hill {shuoxing,tzz}@tamu.edu
@@ 1 Introduction
The recent emergence of powerful Vision Language Models (VLMs) (Li et al., 2022, 2023a; Liu et al., 2024a; Li et al., 2024b; Meta, 2024; Bai et al., 2023; Wang et al., 2024b; Lu et al., 2024; Wu et al., 2024; Bai et al., 2025; Fan et al., 2025) has significantly extended the capabilities of Large Language Models (LLMs) (Devlin et al., 2018; Radford et al., 2019; Brown et al., 2020; Team et al., 2023; Roziere et al., 2023; Touvron et al., 2023a, b; Raffel et al., 2020; Yang et al., 2024; Team, 2024) into the visual domain, paving the way for innovative real-world applications that integrate multimodal information (Moor et al., 2023; Li et al., 2024a; Shao et al., 2024; Xing et al., 2025b; Rana et al., 2023; Kim et al., 2024). Despite their promising performance, VLMs remain susceptible to hallucinations—instances where the model produces outputs containing inaccurate or fabricated details about objects, attributes, and the logical relationships inherent in the input image (Rohrbach et al., 2018; Bai et al., 2024). Several factors contribute to this cross-modal inconsistency, including the separate low-quality or biased training data, imbalanced model architectures, and the disjoint pretraining of the vision encoder and LLM-backbone (Cui et al., 2023; Bai et al., 2024; Zhou et al., 2024a).
To mitigate the hallucinations in VLMs, the Directed Preference Optimization (DPO) techniques have been widely adopted (Deng et al., 2024; Zhou et al., 2024a; Fang et al., 2024; Zhou et al., 2024b; Guo et al., 2024; Chen et al., 2024b; Wang et al., 2024c; Yu et al., 2024b; Li et al., 2023b; Wang et al., 2024a; Xiao et al., 2025; Xie et al., 2024; Fu et al., 2024). This involves constructing datasets enriched with human preference signals specifically targeting hallucinations, and then fine-tuning the models using algorithms like Direct Preference Optimization (DPO) (Rafailov et al., 2024). Existing methods generate the preference data by perturbing the ground truth responses (Zhou et al., 2024a) and corrupting the visual inputs/embeddings (Deng et al., 2024; Amirloo et al., 2024) to generate rejected responses or correcting/refining responses to produce chosen responses (Chen et al., 2024b; Yu et al., 2023a). While methods based on response refinement yield the most reliable preference signals, they face scalability challenges due to the significant costs of manual correction processes. Conversely, directly corrupting input visual information or ground truth responses is overly simplistic, as this brute-force approach fails to generate plausible and natural hallucinations in a controlled manner. Moreover, during fine-tuning, directly applying DPO may cause the model to overly prioritize language-specific preferences, which potentially leads to suboptimal performance and an increased propensity for hallucinations (Wang et al., 2024a).
@@ 2 Preliminaries
Image retrieval aims to find relevant images from large databases – such as vector databases or indexed corpora – based on semantic similarity criteria (Hu et al., 2025). In this paper, we convert all images into vector representations and utilize the cosine similarity metric to evaluate their proximity to a reference image. The similarity between two images, $v_{1}$ and $v_{2}$, is computed as follows:
[…]
where $\cdot$$\cdot$> denotes the inner product in $l_{2}$ space, $f_{p}(v_{i})$ represents the image embeddings generated by the vision encoder $f_{v}(\cdot)$ of VLMs. In this paper, we employ the FAISS library (Douze et al., 2024; Johnson et al., 2019) for efficient vector searches, retrieving the top-$k$ most relevant images.
@@ 3 Methods
As shown in Figure 2, the process begins with an advanced VLM generating chosen responses from input images from the training set. A selective masking process is then applied, strategically omitting segments associated with objects, attributes, or logical relationships identified in the image. Next, leveraging the retrieved image from the same training dataset and the masked responses, the hallucination-prone VLM is prompted to complete the masked elements, obtaining rejected responses. The generated preference pairs (chosen vs. rejected) are then used to fine-tune the VLM with $\mathcal{L}_{\text{rDPO}}$ (eq.˜1), a preference objective that integrates both visual and textual information to penalize hallucinations and reinforce grounded reasoning. Algorithm 1 in Appendix A provides an overview of Re-Align, while the detailed process is explained in the following subsections.
@@ 3.1 Preference Generation
Generating high-quality preference data, which includes both accurate ground-truth responses and controlled hallucinated examples, is crucial for effective preference optimization in pre-trained VLMs. Existing methods construct preference data by perturbing ground-truth responses (Zhou et al., 2024a), corrupting visual inputs/embeddings (Deng et al., 2024; Amirloo et al., 2024) to create rejected responses, or refining responses to obtain chosen responses (Chen et al., 2024b; Yu et al., 2023a). Refinement produces high-quality preference data but comes at a high cost, whereas direct corruption is more scalable yet tends to generate unrealistic hallucinations and fails to produce plausible, natural ones in a controlled manner. To address these limitations, we introduce a novel image retrieval-based pipeline for preference data construction as shown in Figure 3, which consists of three key stages:
@@ 3.2 Preference Optimization
Methods SQA TextVQA $M$-Vet VisWiz LLaVABench MMEP MMEC MMBench Avg. Rank
@@ 4.1 Re-Align for VLMs Alignment
We conducted experiments on both hallucination detection and general VQA tasks. Specifically, we assess our method’s performance in hallucination detection using the POPE dataset (Li et al., 2023c) and HallusionBench (Guan et al., 2023). For general VQA tasks, we leverage a diverse suite of benchmarks including ScienceQA (Lu et al., 2022), TextVQA (Singh et al., 2019), MM-Vet (Yu et al., 2023b), VisWiz (Gurari et al., 2018), LLaVABench (Liu, 2023), MME (Fu et al., 2023), and MMBench (Liu et al., 2024d).
@@
We sample 11k images from the LLaVA-Instruct-150K dataset (Liu et al., 2024a) to construct preference data, as illustrated in Figure 3. These images are initially used to generate QA pairs based on image captions and simple VQA tasks using GPT-4o mini (OpenAI, 2024). Furthermore, the images are encoded using clip-vit-large-patch14 (Radford et al., 2021a) to construct the knowledge base for image retrieval. For rejected responses, we use GPT-4o mini to mask the chosen response and all-mpnet-base-v2 (Reimers and Gurevych, 2019) to compute the similarity between the completed masked response and the original chosen response. We use LLaVA-v1.5-7B (Liu et al., 2024a) and LLaVA-v1.6-Mistral-7B (Li et al., 2024b) as our backbone models and perform Re-Align fine-tuning for $1$ epoch. All evaluations are conducted with a temperature setting of 0, and baseline results are reproduced using open-sourced model weights.
@@ 4.2 Scalability and Generalizability
  • Image-to-Text models: the typical architecture of VLMs, where a vision encoder is integrated with an LLM to enable cross-modal understanding. In this section, we evaluate Re-Align on LLaVA-v1.5-7B (Liu et al., 2024a), LLaVA-v1.5-13B (Liu et al., 2024a), LLaVA-v1.6-Vicuna-7B (Li et al., 2024b), LLaVA-v1.6-Vicuna-13B (Li et al., 2024b), Qwen2.5-VL-3B-Instruct (Bai et al., 2025), and Qwen2.5-VL-7B-Instruct (Bai et al., 2025).
@@
Qwen2.5-VL-
    3B-Instruct
88.32 87.60 86.63
w. Re-Align 89.69↑1.37 88.33↑0.73 87.16↑0.53
Qwen2.5-VL-
    7B-Instruct
88.73 87.90 86.87
w. Re-Align 89.27↑0.54 88.10↑0.20 87.10↑0.23
@@ Results
Table 3 presents the performance of Re-Align using both standard image-to-text and unified VLM backbones across model sizes from 1B to 13B on the POPE benchmark (Li et al., 2023c). In experiments with the LLaVA-v1.5 series (Liu et al., 2024a), none of the baseline approaches consistently improve performance for either the 7B or the 13B models, highlighting the limited scalability of these methods. In contrast, Re-Align achieved substantial performance gains, outperforming both the baseline models and the vanilla version—most notably on the LLaVA-v1.5-13B variant. Similarly, experiments with the LLaVA-v1.6-Vicuna series (Li et al., 2024b) and Qwen2.5-VL series (Bai et al., 2025) revealed the same trend, further underscoring Re-Align’s superior scalability. For unified vision-language models, especially Janus-Pro, integrating Re-Align yields a significant performance boost. Notably, Janus-Pro-1B experiences the greatest improvement, underscoring Re-Align’s robustness across different model architectures. However, Janus-Pro-1B, being the smallest among the evaluated VLMs, also exhibits the poorest overall performance on POPE, suggesting a correlation between model size and the propensity for hallucinations.
@@ 4.3 Ablation Study
In this section, we conduct a comprehensive ablation study to explore how the data curation framework and design of the objective function affect the Re-Align’ performance. The experimental setup follows the same setting as VLMs alignment experiments, with LLaVA-1.5-7B as the backbone.
@@
Table 4: Impact of similarity threshold $\tau$ for generating the rejected responses in Re-Align across general and hallucination benchmarks for VLMs.
@@
In data curation, we generate preference data by inducing hallucinations at the segment level. To further investigate the impact of finer-grained perturbations, we conduct experiments using sentence-level masking. As shown in Table 5, using a sentence-level masking strategy, Re-Align still demonstrates significant improvement in reducing hallucination in VLMs. However, this approach leads to a slight drop in performance on general VQA tasks. More discussions on the masking strategy can be found in Appendix 5.
Masking
Strategy
SQA TextVQA POPEr POPEp POPEa
@@
Table 5: Impact of masking strategy across general and hallucination benchmarks for VLMs.
@@
Training Epochs
For a fair comparison with prior baselines, we primarily report results of Re-Align under a one-epoch fine-tuning setup, which already demonstrates the effectiveness of our proposed method. To further explore the impact of training duration, we conduct additional experiments with extended fine-tuning of up to three epochs.
Num
Epoch
SQA TextVQA POPEr POPEp POPEa
1 68.10 58.55 88.65 87.43 85.16
2 68.27 58.47 88.91 87.52 85.16
3 68.17 58.60 88.57 87.60 85.43
Table 7: Impact of the number of training epochs across general and hallucination benchmarks for VLMs.
As shown in Table 7, Re-Align exhibits stable performance across longer training schedules, with results consistently maintained and in some cases slightly improved on both general VQA benchmarks (SQA, TextVQA) and hallucination benchmarks (POPE). This indicates that our method is robust to extended training and not prone to overfitting, while continuing to deliver reliable gains.
@@ 5 Discussions
Performance Variations on General VQA tasks
Segment-level Preference
Building on the findings of Yu et al. (2024b), we generate preference data by inducing hallucinations at the segment level rather than at the sentence level (as seen in approaches such as POVID (Zhou et al., 2024a), STIC (Deng et al., 2024), and CSR (Zhou et al., 2024b)), to provide robust supervision signals during the alignment process. This finer-grained preference modeling yields clearer and more precise learning signals, enabling the model to better distinguish between subtle hallucinations and ground truth responses. To further investigate these segment-level preference signals, we expanded the fine-tuning dataset from $11k$ to $16k$ image samples. As illustrated in Figure 4, when using LLaVA-v1.6-Mistral-7B as the backbone with ScienceQA as the case study, Re-Align achieved a significant performance improvement—from $0.45$ to 1.341.34—demonstrating the effectiveness of our approach.
Computational Complexity
The proposed Re-Align pipeline can be modularized into offline preprocessing and online training integration (detailed computational cost can be found in the Appendix):
  • Preprocessing: Image retrieval, strategic masking, and preference pair generation can be entirely performed offline as a one-time data preprocessing step. This includes CLIP-based similarity search, mask generation, and SentenceTransformer-based similarity computation. Once completed, these preprocessed preference pairs can be reused across multiple training runs without additional overhead.
  • Training Overhead: The actual training process introduces minimal additional computational overhead ( 5-10% increased training time) compared to standard DPO, with virtually identical memory requirements. The additional cost stems only from:
    • Forward passes through the visual encoder for retrieved images;
    • Generation passes through the LLM backbone for computing the vDPO loss component.
@@ 6 Related Work
Reinforcement Learning from Human Feedback (RLHF) has emerged as a crucial technique for incorporating human preference signals into machine learning methods and models (Dong et al., 2024; Yin et al., 2022). RLHF frameworks can be broadly categorized into deep RL-based approaches and direct preference learning approaches. In deep RL-based methods, a reward model is first constructed, after which Proximal Policy Optimization (PPO) (Schulman et al., 2017; Christiano et al., 2017; Ziegler et al., 2019) is employed to optimize the reward signals with KL regularization (Ouyang et al., 2022; Touvron et al., 2023b). While the direct preference learning approaches optimize a designed loss target on the offline preference dataset directly, eliminating the need for a separate reward model (Rafailov et al., 2024; Azar et al., 2024; Tang et al., 2024; Ethayarajh et al., 2024).
@@
Large Vision Language Models (VLMs) (Li et al., 2022, 2023a; Liu et al., 2024a; Li et al., 2024b; Meta, 2024; Bai et al., 2023; Wang et al., 2024b; Lu et al., 2024; Wu et al., 2024; Bai et al., 2025; Fan et al., 2025; Abouelenin et al., 2025) extended the understanding and reasoning capabilities of Large Language Models (LLMs) (Devlin et al., 2018; Radford et al., 2019; Brown et al., 2020; Team et al., 2023; Roziere et al., 2023; Touvron et al., 2023a, b; Raffel et al., 2020; Yang et al., 2024; Team, 2024; Pan et al., 2024; Yang et al., 2025) into the visual domain. By integrating vision encoders, such as CLIP (Radford et al., 2021b), image patches are first converted into embeddings and then projected to align with text embedding space, unlocking unprecedented cross-modal applications in the real world, such as biomedical imaging (Moor et al., 2023; Li et al., 2024a; Zuo et al., 2025), autonomous systems (Shao et al., 2024; Tian et al., 2024; Sima et al., 2023; Xing et al., 2025b; Ma et al., 2025; Wang et al., 2025b; Li et al., 2025b; Gao et al., 2025b), and robotics (Rana et al., 2023; Kim et al., 2024; Xing et al., 2025c).
@@
Current VLMs often suffer from hallucinations, producing inaccurate or misleading information that fails to accurately represent the content of the provided image (Zhu et al., 2024; Bai et al., 2024; Qian et al., 2025; Xing et al., 2025a). Such misalignments can have catastrophic consequences when these models are deployed in real-world scenarios (Xing et al., 2024). To address cross-modality hallucinations, recent research has primarily focused on applying direct preference optimization (Deng et al., 2024; Zhou et al., 2024a; Fang et al., 2024; Zhou et al., 2024b; Guo et al., 2024; Chen et al., 2024b; Wang et al., 2024c; Yu et al., 2024b; Li et al., 2023b; Wang et al., 2024a) or contrastive learning (Sarkar et al., 2024) on the curated datasets with preference signals, and utilizing model editing techniques (Liu et al., 2024b; Yu et al., 2024a).